Please Help How do you enter a formula into the computer using exel (graphing) so the computer will understand it? Regarding algebra thank you
Answer by adamchapman(301)

The way I do it is by making a column of x and a column of y.
Put the value "0" in the cell A1 (just type the zero key into the cell) now decide what increments you want your x-values to have. Imagine we wanted to plot x in incremtnts of 0.2.
We would type:
=A1+0.2
in the cell A2 and press enter.
Now click on the cell A2 and drag downwards. This will add 0.2 to the cell above for all the highlighted cells.
Now the y colmn:
Imagine we wanted to graph y=3(x^2).
We would type:
=3*(A1^2)
into the cell B1 and press the enter key. You MUST notice to use a star (*) to multiply. Now click and drag the cell B2 downwards to match y values with all the listed values of x.
Now highlight all your cells with values in them in columns A and B.
Click the "Insert" menu at the top of the page.
Click on "Chart"
The chart wizard will pop-up. The wizard explains the rest of the process as you go, but if you need any more help please don't hesitate to contact me at adam.chapman@student.manchester.ac.uk
Adam